OpenStudy (anonymous):

Which of the following inequalities matches the graph? -6x + y < 3 6x + y < 3 6x - y < -3 The correct inequality is not listed

OpenStudy (anonymous):

where is the graph?

OpenStudy (ivettef365):

By looking at the graph, you have that the y-int is (0,3)

OpenStudy (anonymous):

@SavannahSaturn

OpenStudy (ivettef365):

The slope is a positive slope

OpenStudy (ivettef365):

Then convert to formula y=mx+b

OpenStudy (ivettef365):

Where m = slope and b is the y-int

OpenStudy (anonymous):

okay the graph is up there @T.E.D..

OpenStudy (anonymous):

I still don't get it

OpenStudy (ivettef365):

The only possible answer would be -6x +y <3

OpenStudy (ivettef365):

Because when you convert as y=mx+b then you have y <6x+3 slope = 6 (positive0 y-int = 3